OpenStudy (boonerileycat):

6. Given the following equation. How much heat is released by the decomposition of 36g of water. 2H2O + 25kJ  2H2 + O2

OpenStudy (photon336):

This is your reaction right? \[2H_2O + 25kj \rightarrow 2H_2 + O_2\]

OpenStudy (photon336):

you mean heat is absorbed? because heat is on the reactant side which means that the decomposition of water is and endothermic process.

OpenStudy (boonerileycat):

Yes that is the correct equation and I would suppose what you said was probably right! @Photon336

OpenStudy (photon336):

yeah, the decomposition of water is endothermic and requires the input of 285.8 kJ energy per mole of water decomposed. so what you would need to do is to find out how many moles of water are in 36 grams of water. just curious where did you get 25KJ from?

OpenStudy (boonerileycat):

From the sheet my Chem teach posted. I have no idea were he got it from but this was the only equation I was stuck on. Thank you for the help! Could you explain how to find the 36 grams of water? @Photon336

OpenStudy (photon336):

So one mole of water \[\frac{ mole~H_2O }{ 18~grams }\] we figure out the molar mass of water which we go to the periodic table and add up the molecular masses of each. for Hydrogen its 1 and for oxygen its 16 we have 2 hydrogens, so we multiply the number of atoms by molar mass. since we have 2 hydrogens it's 2*1 = 2 and for oxygen it's just 16 because we've got 1 oxygen atom. now we add them all up together 2+16 = 18 grams so to figure out the number of moles in 36 grams we do this we multiply the number of grams by the molar mass. you notice something? the grams cancel out and we're left with moles. \[36~grams*(\frac{ mole~H_2O }{ 18~grams }) = \frac{ 36~moles }{ 18 } = 2~moles~H2O\] Well, I guess if that's the number your teacher gave you, you should probably use that number.

OpenStudy (photon336):

well, the thing is usually numbers like 25 kJ are reported in kJ/moles

OpenStudy (boonerileycat):

That's the answer? @Photon336

OpenStudy (photon336):

Well, the thing is do you understand how I got 2 moles of H2O ?

OpenStudy (photon336):

so i think that it would mean to decompose 2 moles of H2O, you would need to put in 25kj of energy.

OpenStudy (boonerileycat):

@Photon336 I got it right!! And I understood, thank you so much